OpenStudy (hockeychick23):

Suppose a normal distribution has a mean of 38 and a standard deviation of 2. What is the probability that a data value is between 36 and 43?

OpenStudy (anonymous):

U should ask hartnn this he is the best with math

OpenStudy (calculusfunctions):

First step is to determine the z-score for each data. Then look at your table that provides the probabilities of the z-scores to find the corresponding probabilities. Then subtract the probability of the smaller z-score from the probability of the larger z-score.

OpenStudy (hockeychick23):

ok so i know that the z-score= x-myew/standard deviation, but do you know what the x would be? because then i get: z-score= x-38/2

OpenStudy (calculusfunctions):

You are given the x's in the question (36 and 43).

OpenStudy (hockeychick23):

ohh so i use both separately ok thanks!

OpenStudy (calculusfunctions):

The x's are your data values.

OpenStudy (hockeychick23):

so now i have 36-38/2= -1 and 43-38/2=5/2

OpenStudy (calculusfunctions):

Yes, you have to find the z-score for each data value (x). Follow the steps I laid out above.

OpenStudy (hockeychick23):

ok, i found the z-score for -1= .1587 and z-score for 2.5=.9938 .9938- .1587= .8351

OpenStudy (calculusfunctions):

I don't have the table in front of me so I won't be able to verify your final answer, however if you follow the steps and perform the calculations and use the table properly, then you should be fine.

OpenStudy (hockeychick23):

ok thanks!

OpenStudy (calculusfunctions):

Alright then GOOD LUCK!!